Aplane travels 250 mph in still air. The plane goes 560 miles with wind and 440 miles against wind in the same time. What is speed of wind?
Answer by lwsshak3(11628)

Aplane travels 250 mph in still air. The plane goes 560 miles with wind and 440 miles against wind in the same time. What is speed of wind?
**
let w=speed of wind
250+w=speed of plane downwind
250-w=speed of plane against wind
Travel time=distance/speed
..
560/(250+w)=440/(250-w)
560*250-560w=440*250+440w
560*250-440*250=440w+560w
1000w=250(560-440)=250*120=30000
w=30
speed of wind=30 mph